HOW MANY FOUR DIGIT NUMBERS CAN BE FORMED FROM 1,2,3,4,5,6,7 WHICH IS GREATER THAN 3400?

Answer by sudhanshu_kmr(1152)


No. of 4 digit numbers where first digit is 4,5,6 or 7.
4* 6P3 = 4 * 120 = 480
No. of 4 digit numbers where first digit is 3..
second place can be filled by 4,5,6,or 7.
no. of possible numbers = 4*5P2 = 4 * 20 = 80

total no. of possible numbers formed from 1,2,3,4,5,6 and 7
and greater than 3400 = 480 + 80 = 560

Note- repetition is not allowed here.